Consider an economy with the following data:

Consumption expenditure = $271,650.0 million
Planned investment = $218,972.0 million
Government expenditure = $58,666.0 million
Export expenditure = $940,148.0 million
Import expenditure = $820,652.0 million
Autonomous taxes = $306,700.0 million
Income tax rate = 20%
Marginal propensity to save = 0.6
Marginal propensity to import = 0.1

Calculate the level of total savings when the level of income equals $620,240 million (solve to one decimal point).
Answer #1

Answer

Total Saving = Private Saving + Public Saving

Private Saving = Y - C - T

where Y = Income =  $620,240 million

C = Consumption Expenditure = $271,650.0 million

T = Total Taxes

=> Private Saving = Y - C - T = 620,240 million - 271,650.0 million - T

Public Saving = T - G

where G = Government Purchases = $58,666.0 million

Public Saving = T - G = T - 58,666.0 million

Hence, Total Saving = Private Saving + Public Saving

=> Total Saving = 620,240 million - 271,650.0 million - T + T - 58,666.0 million

Hence, Total Saving = $289924.0 million.